A ball is projected vertically up from ground.Boy 'A' Standing at the window of first floor of a nearby building observes that the time interval between the ball crossing him while going up and the ball crossing him while going down is 2 s. Another boy 'B' standing on the second floor notices that time interval between the ball passing him twice during up motion and down motion is 1 s.Calculate the difference between the vertical positions of boy B and boy A.
(Assume g=10m/s^(2))
Ans: 3.75m
1
Expert's answer
2020-01-03T09:18:41-0500

Let "v_0" is an initial speed of the ball. The total time of motion of the ball is 2 sec, so initial velocity


"v_0=gt/2=10\times 2/2=10\:\rm m/s."

The time of motion of the ball between two boys is "\tau=0.5" sec. Hence the difference between the vertical positions of boy B and boy A


"d=v_0\tau-g\tau^2/2""=10\times 0.5-10\times 0.5^2/2=3.75\:\rm m"